Portfolio Weights An investor owns $23,000 of Adobe Systems stock, $26,000 of Dow Chemical, and $36,000 of Office Depot. What are the portfolio weights of each stock?

Multiple Choice

  • Adobe System = 0.27, Dow Chemical = 0.31, Office Depot = 0.42

  • Adobe System = 0.27, Dow Chemical = 0.33, Office Depot = 0.40

  • Adobe System = 0.33, Dow Chemical = 0.33, Office Depot = 0.33

  • Adobe System = 0.31, Dow Chemical = 0.27, Office Depot = 0.42
Answer #1

Weight of Adobe System = 23,000/(23,000 + 26,000 + 36,000)

Weight of Adobe System = 0.27

Weight of Dow Chemical = 26,000/(23,000 + 26,000 + 36,000)

Weight of Dow Chemical = 0.31

Weight of Office Depot = 36,000/(23,000 + 26,000 + 36,000)

Weight of Office Depot = 0.42

Adobe System = 0.27, Dow Chemical = 0.31, Office Depot = 0.42